# The evolution of trust as a cognitive shortcut in repeated interactions

### v2

@@ 3 Results
We begin our analysis by clarifying the role of reciprocal strategies in promoting cooperation across the range of pairwise social dilemmas defined by the Prisoner’s Dilemma, Snowdrift and Stag-Hunt games. We then determine how reciprocal strategies are affected by the realistic assumption of an opportunity cost of verifying the partner agent’s actions, $\epsilon>0$, and how the effect of the opportunity cost differs between games. From this, we can then proceed to analyse the evolutionary advantage of trust, and its effectiveness in promoting cooperation across the social dilemma games. We initially assume, unless otherwise specified, that $\mu_{\mathrm{e}}=0$, i.e. agents do not make errors in their actions, unless otherwise specified.
@@ 3.1 When does reciprocity promote cooperation in different social dilemmas?
Figure 2: Frequency of the Tit-for-Tat (TFT) strategy (left), and the impact on the frequency of cooperation of adding TFT to the pool of available strategies AllC and AllD(right). The results are shown as a function of temptation to defect $T$ and sucker’s payoff $S$, and in the absence ($\epsilon=0$, top panels) or presence ($\epsilon=0.25$, bottom panels) of an opportunity cost of verifying the partner’s actions. Prisoner’s Dilemma (PD) games occur in the lower right quadrant of $S$-$T$ space, Snowdrift (SD) games occur in the top right quadrant and Stag-Hunt (SH) games in the lower left quadrant. The red dots show the particular game instances that we use for later analysis. Parameters: a game with $r=50$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
@@ 3.1.1 The role of reciprocity across social dilemma games
We first examine how Tit-for-Tat (TFT) performs across the space of Prisoner’s Dilemma (PD), Snowdrift (SD), and Stag-Hunt (SH) games, and compute how much introducing TFT into a strategy pool that initially consists only of Always Cooperate (AllC) and Always Defect (AllD) promotes cooperation. We start by considering the case where there are no opportunity costs to verifying the partner agent’s actions. Our results, presented in the top part of Figure 2, show that TFT is a frequent strategy across all the social dilemma games, and is more frequent when the temptation to defect, $T$, is high. TFT is at the highest frequency in PD types of interaction, and less frequent in SD and in SH situations. The top right panel in Figure 2, indicates that adding TFT into the pool of available strategies also increases the frequency of cooperation across all three social dilemma games.
3.1.2 Tit-for-Tat and the opportunity cost of verifying a partner’s actions
Conducting the same analysis, but this time in the presence of an opportunity cost ($\epsilon=0.25$), our results (bottom panels of Figure 2) show that including the opportunity cost drastically reduces the frequency of TFT (compare the top left and bottom left panels). The drop in frequency is the strongest in SH situations, where TFT is close to absent for many SH games in $S$-$T$ space. By contrast, the drop in frequency of TFT is smallest for PD games. Opportunity costs also reduce the effect of the presence of TFT on increasing cooperation across all games. This effect is less pronounced in SH games, even though the frequency of TFT is lowest in this part of $S$-$T$ space compared to the other games. Thus, even a low frequency of TFT promotes cooperation in the risky coordination situations represented by the SH game.
@@ 3.1.2 Tit-for-Tat and the opportunity cost of verifying a partner’s actions
Figure 3 examines which strategies foster cooperation in different games, and demonstrates that in the PD, cooperation primarily arises from the high frequency of TFT. In contrast, in the SH and SD games, cooperation is due to a mix of AllC and TFT strategies. When considering opportunity costs of verifying the partner agent’s actions, this distinction becomes even more pronounced, with nearly all cooperation in the SH game resulting from the higher frequency of AllC. This finding explains why opportunity costs have high impacts in the PD, where cooperation results directly from TFT individuals cooperating with each other. But there is less impact in the SH, where cooperation results from AllC players that are immune to these costs since their behaviour is unconditional on their partner’s actions. Now that we have established how TFT fares across different social dilemmas, and how a cost of verifying partner’s action affects this, we now conduct the same analysis for trust-based strategies.
@@
Figure 3: Frequency of cooperators using the Tit-for-Tat strategy, in the absence of the opportunity cost of verifying the partner’s actions (top), and with the cost $\epsilon=0.25$ (bottom). Parameters: a game with $r=50$, and evolutionary dynamics where $\beta=0.1$ and $N=100$.
3.2 Trust across games
@@
We now investigate, as a function of the opportunity cost, (i) the frequency of the Trust-based Cooperation strategy TUC when all five strategies are possible, and (ii) the impact of adding trust-based strategies (TUC and TUD) to the pool of available strategies on the frequency of cooperation.
@@
Figure 4: Left: Frequency of strategies as a function of the opportunity cost, $\epsilon$. Right: Frequency of cooperation in the absence or presence of trust-based strategies TUC and TUD, as a function of the opportunity cost, $\epsilon$. The difference in frequency of cooperation when trust-based strategies are included is shaded in green where positive and red where negative. Results are presented for one example of each of the Prisoner’s Dilemma, Snowdrift and Stag-Hunt games. Parameters: trust-based strategies with $\theta=3$ and $p=0.25$, games with $r=50$, and evolutionary dynamics where $\beta=0.1$ and $N=100$.
@@
Figure 5: Frequency of the TUC strategy (left) and the impact of considering trust strategies on the frequency of cooperation (right), as a function of the temptation to defect $T$ and the sucker’s payoff $S$. The difference in frequency of cooperation is calculated between a system with all five strategies and a system with only AllC, AllD and TFT. The red dots correspond to the values of S and T used in Figure 4. Parameters: trust-based strategies with $\theta=3$ and $p=0.25$, games with $\epsilon=0.25$ and $r=50$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
@@
3.2.1 Baseline analysis: Trust is a successful strategy and promotes cooperation in the repeated Prisoner’s Dilemma
@@
In the top part of Figure 4, which focuses on the PD, our analysis replicates previous findings [24]. First, increasing opportunity costs diminish the frequency of the TFT strategy (top left panel), since TFT must pay the cost of verifying every round. Second, higher opportunity costs favour TUC compared to TFT, with TUC being the most successful strategy overall for intermediate opportunity costs in the PD. Third, the inclusion of trust-based strategies (TUC and TUD) in a system that previously included only AllC, AllD and TFT leads to an overall increase in the frequency of cooperation (top right panel). When opportunity costs are high, the AllD strategy becomes the most prevalent, but (i) TUC remains among the most common across a wide range of opportunity costs, and (ii) the inclusion of trust-based strategies always (for the studied range of opportunity costs) increases the frequency of cooperation.
@@
3.2.2 Trust is a frequent strategy across different games
@@
Our setup now allows us to examine if these findings generalise across different types of games, and thus different types of social interaction. Our results in Figure 4 and the left panel of Figure 5 demonstrate that Trust-based Cooperation (TUC) consistently remains a frequent strategy across the three different games, and for a broad range of opportunity costs. In other words, trust is expected to be a widespread heuristic across various types of interaction, not just in PD situations. A notable difference between the games is that trust-based strategies are highly prevalent in the PD, while AllC is almost absent. In this situation, trust is a successful strategy, but naive unconditional cooperation is not. For the SD and SH, both TUC and AllC strategies are frequent, indicating that a mix of careful trusters and naive unconditional cooperators could be observed in these more relaxed social dilemmas.
@@
3.2.3 Trust increases the frequency of cooperation when the temptation to defect is high
@@
Our results reveal that the impact of incorporating a trust-based strategy on cooperation differs across the various games. Specifically, while trust is always beneficial in the PD, it has a limited effect on the frequency of cooperation in SD games and in SH games. As shown in Figure 5, trust-based strategies generally increase cooperation when the temptation to defect (T) is high. However, trust does not seem to increase cooperation in SH coordination games, even though it is present at a significant frequency. This is because in coordination games, we have shown earlier that reciprocal strategies are only a springboard for individuals that always cooperate to get in place. Once in place, these unconditional strategies can enjoy a high payoff without ever paying opportunity costs.
3.3 Trust when individuals make errors
@@
Finally, we consider cases where agents can make unintentional errors in their actions. Figure 6 shows the frequency of the different strategies and of cooperation as a function of the error rate, $\mu_{\mathrm{e}}$. Our results show that an increasing error rate decreases the frequency of TUC across games, in favour of the strategy AllD. However, our results also show that the frequency of TUC remains considerable even for quite large error rates such as 5%.
@@
Figure 6: Frequency of strategies (left) and frequency of cooperation in the absence or presence of trust-based strategies TUC and TUD (right), as a function of the error rate, $\mu_{\mathrm{e}}$. For clarity, the difference in frequency of cooperation is shaded in green when positive and red when negative. Parameters: trust-based strategies with $\theta=3$ and $p=0.25$, games with $\epsilon=0.25$ and $r=50$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
@@
The right hand side of Figure 6 shows that including error rates up to 10% does not lessen the beneficial effect of trust on the frequency of cooperation in all of the games, and can even increase it. For instance, TUC increases cooperation the most in the PD when the error rate is around 2%-3%. Another example is that trust-based strategies improve cooperation when the error rate is above 0 in the SH, in contrast to the results where agents do not make errors.
@@
In other words, the presence of errors, if not too frequent, can still select for trust-based strategies and make them even more beneficial for cooperation. Figure 7 shows that these conclusions hold for games repeated for a larger number of rounds ($r=100$), and where trust-based strategies verify less frequently once the trust threshold has been reached (high trustfulness, $p=0.04$).
@@
The effect of high trustfulness on cooperation varies with the type of game. Specifically, lowering $p$ from 0.25 to 0.04 increases cooperation in SD and SH games, but it decreases cooperation in the PD (right-hand pane of Figure 7). Thus, the optimum level of trustfulness for promoting cooperation depends on the type of social dilemma, with more relaxed dilemmas favouring greater trustfulness.
@@
Figure 7: Difference in the frequency of cooperation when trust-based strategies TUC and TUD are included, as a function of the error rate, $\mu_{\mathrm{e}}$, for three different social dilemmas and for long interactions or high trustfulness. Parameters: trust-based strategies with $\theta=3$, games with $\epsilon=0.25$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
@@ 4 Discussion
Our results can be summarised in three points. First, we have shown that trust, as a heuristic to avoid the opportunity cost of verifying its partner’s action, is a successful strategy that can evolve in a wide range of interactions. Importantly, this happens even in the presence of exploitative strategies (TUD) tailored to take advantage of this trust.
@@
Second, by disentangling behavioural measures of trust from behavioural measures of cooperation, we have been able to rigorously examine the effect that trust has on the overall level of cooperation in a population. Our results show that the presence of the Trust-based Cooperation strategy (TUC) increases the overall level of cooperation when the temptation to defect is high, but not in coordination dilemmas such as Stag-Hunt games. We have shown that this is because reciprocity in coordination games is useful to kickstart cooperation, but is not necessary once cooperation is established. This suggests that interventions aiming to promote the success of TUC, such as reducing high opportunity costs, would have less effect on cooperation in risky coordination games compared to stricter social dilemmas.
@@
Third, we have shown that when there is the possibility of agents making errors in their actions, then TUC remains frequent, and can even be more beneficial for cooperation compared to the case where agents never make errors. Notably, TUC does not suffer as much from errors as TFT, as TUC agents ignore errors from cooperative partners on the rounds that they do not verify. By contrast, TFT is not trustful and verifies its partner’s action every round, which means that one single mistake causes TFT to defect in retaliation. This can then cause the partner to defect in retaliation, leading to a loss of cooperation over many rounds resulting from one mistake [38]. Trust can prevent this spiral into defection.
@@
Our model is based on the assumption that it can be costly to know whether the action of another agent was actually cooperative – even when you observe that action [24]. Everyday examples of this are abundant. For example, there is a time cost to verifying if the correct amount has been debited from your bank account, or if goods that you have ordered online are genuine rather than fake. When interacting with artificial intelligence (AI), these costs can become even greater. For example, are personalised music recommendations from a streaming app influenced by paid promotions [39]? Has an LLM agent completed a task in a way that meets the preferences I prompted it with, or were its actions not in my interests due to biases in its training data [40]?
@@
Our operationalisation of trust – as choosing not to verify another agent’s actions – provides an objective behavioural measure that applies across a broad spectrum of social interactions. Using this, trust can be measured in experimental settings by how often someone chooses (or not) to pay a cost to verify another agent’s actions. This could be carried out in standard behavioural economics settings [41], where a participant is given the option to pay to find out the action of its partner in each round of a repeated social dilemma game. Some experimental studies have already considered costly verification for the particular case of the “trust game” [42, 43]. In this game, one player is the investor and the other player is the trustee. The investor begins with an endowment amount of money. They then choose a proportion (between 0 and 1 inclusive) of this to send to the trustee. The amount sent is multiplied by a factor >1>1. The trustee then chooses the proportion of this to return to the investor. When the game is repeated, most experiments assume that the investor is given complete information about the previous actions of the trustee at no cost. However, a study has shown that when a cost to obtain this information is introduced, people are sensitive to this cost and verify less [42]. Moreover, the frequency of verifying was lower in later rounds, in a pattern akin to the TUC strategy. Future work should extend these experiments beyond the trust game and into the larger space of social dilemma games that we have modelled.
@@
Crucially, this trust measure can also be applied in more realistic experimental settings. One important application of this is experiments to measure the factors affecting people’s trust in AI. Although there have been many empirical studies in computer science that have aimed to measure factors affecting trust in various applications of AI, these studies have lacked a unified theoretical framework [44, 45]. This has led to conflicting results and a lack of predictive models, with “Trustworthiness mechanisms and measures […] being advanced in AI regulations and standards that may not actually increase trust” [46]. The trust measure we have introduced here can help to resolve this impasse, by providing an objective measure of trust that is grounded in (evolutionary) game theory. For example, incentivised experiments could be designed where a user can take advice from an AI agent in a realistic setting, such as an e-commerce chatbot recommending products or an LLM answering questions on a text, and can pay a cost to verify whether that advice is correct. This can then be connected with EGT models by varying factors such as how frequently the AI gives correct information (i.e. its strategy to cooperate or defect), or whether the AI has a high reputation from other users.
@@
The costs of verifying also lead to problems for the alignment and control of AI, that is, ensuring that the goals and behaviours of AI systems are aligned with human values and intentions. Research in AI safety has stressed the risk of scheming – AI systems that deceptively appear to be aligned with human goals during training, but then pursue misaligned goals after deployment once they are no longer closely monitored [47]. This type of scheming corresponds to the TUD strategy in our model – cooperate until an auditor or developer relaxes oversight, and then defect. The opportunity cost of verifying maps to the cost of auditing the AI system. Extensions of our model, with the payoff matrix calibrated to the parameters used in red-team blue-team AI control experiments [48], could therefore provide insights into auditing regimes and other institutional incentives that minimise vulnerability to scheming by LLMs and other future AI systems.
@@
Most generally, our model demonstrates the evolutionary success of a trust heuristic in competition with strategies that do not use trust (ALLC, ALLD and TFT), and in competition with strategies designed to build and then exploit trust (TUD). Trust evolves even though trust-based strategies are costly to use compared to unconditional strategies (since they pay the cost of verifying before trust is established and periodically thereafter), and are risky compared to standard reciprocal strategies (since they could potentially be exploited once they stop verifying every round). These results suggest that a trust heuristic could have evolved in humans by natural selection. Indeed, there is evidence that human trust behaviour is to some extent genetically determined [49], and hence is an evolved trait. Our results suggest that repeated social dilemma games, as played by humans throughout (pre)history [50], could provide selection pressure for the evolution of trust.
@@
There are several limits to our model which could be addressed in future work. First, we have not modelled the evolution of the parameters of trust-based strategies that control when they verify their co-player’s actions, $\theta$ and $p$. Future work could allow these to be evolving traits, allowing evolutionary optimal values to be found for different game parameters.

### v3

@@ 3 Results
We begin our analysis by clarifying the role of reciprocal strategies in promoting cooperation across the range of pairwise social dilemmas defined by the Prisoner’s Dilemma, the Snowdrift, and the Stag-Hunt games. We then determine how reciprocal strategies are affected by the realistic assumption of a cost of monitoring a partner’s actions, $\varepsilon>0$, and how the effect of the monitoring cost differs between games. From this, we proceed to analyse the evolutionary advantage of trust-based (reduced monitoring) strategies, and their effectiveness in promoting cooperation across the different social dilemma games. We initially assume, unless otherwise specified, that $\mu_{\mathrm{e}}=0$, i.e. agents do not make errors in their actions, before relaxing this assumption in Section 3.3.
@@ 3.1 When does reciprocity promote cooperation in different social dilemmas?
Figure 2: Frequency of the Tit-for-Tat (TFT) strategy (left), and the impact on the frequency of cooperation of adding TFT to the pool of available strategies AllC and AllD (right). The results are shown as a function of the temptation to defect $T$ and sucker’s payoff $S$, and in the absence ($\varepsilon=0$, top panels) or presence ($\varepsilon=0.25$, bottom panels) of the cost of monitoring a partner’s actions. Prisoner’s Dilemma (PD) games occur in the lower right quadrant of $S$-$T$ space, Snowdrift (SD) games occur in the top right quadrant and Stag-Hunt (SH) games occur in the lower left quadrant. The red dots show the particular game instances that we use for later analysis. Parameters: a game with $r=50$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
@@ 3.1.1 The role of reciprocity across social dilemma games
We first examine how Tit-for-Tat (TFT) performs across the space of Prisoner’s Dilemma (PD), Snowdrift (SD), and Stag-Hunt (SH) games, and compute how much introducing TFT into a strategy pool that initially consists only of Always Cooperate (AllC) and Always Defect (AllD) promotes cooperation. We start by considering the case where there is no cost of monitoring a partner’s actions. Our results, presented in the top part of Figure 2, show that TFT is a frequent strategy across all social dilemma games, and is more frequent when the temptation to defect, $T$, is high. TFT is at the highest frequency in PD games, and less frequent in SD and in SH games. The top right panel in Figure 2 indicates that adding TFT into the pool of available strategies also increases the frequency of cooperation across all three social dilemma games.
3.1.2 Tit-for-Tat and the cost of monitoring a partner’s actions
Conducting the same analysis, but this time in the presence of a monitoring cost ($\varepsilon=0.25$), our results (bottom panels of Figure 2) show that including a monitoring cost drastically reduces the frequency of TFT (compare the top left and bottom left panels). The drop in frequency is largest in SH situations, where TFT is close to absent for many SH games in $S$-$T$ space. By contrast, the drop in frequency of TFT is smallest for PD games. A monitoring cost also reduces the effect of the presence of TFT on increasing cooperation across all games. This effect is less pronounced in SH games, even though the frequency of TFT is lowest in this part of $S$-$T$ space compared to the other games. Thus, even a low frequency of TFT promotes cooperation in the risky coordination situations represented by the SH game.
@@ 3.1.2 Tit-for-Tat and the cost of monitoring a partner’s actions
Figure 3 examines which strategies foster cooperation in different games, and demonstrates that in the PD, cooperation primarily arises from the high frequency of TFT. In contrast, in the SH and SD games, cooperation is due to a mix of AllC and TFT strategies. When considering costs of monitoring the partner’s actions, this distinction becomes even more pronounced, with nearly all cooperation in the SH game resulting from the higher frequency of AllC. This finding explains why a monitoring cost has a strong impact in the PD, where cooperation results directly from TFT individuals cooperating with each other. But there is less impact in the SH, where cooperation results from AllC players that are immune to these costs since their behaviour is unconditional on their partner’s actions. Now that we have established how TFT fares across different social dilemmas, and how a cost of monitoring a partner’s actions affects this, we conduct the same analysis for trust-based strategies.
@@
Figure 3: Frequency of cooperators using the Tit-for-Tat strategy, in the absence of the cost of monitoring a partner’s actions (top), and with the cost $\varepsilon=0.25$ (bottom). Parameters: a game with $r=50$, and evolutionary dynamics where $\beta=0.1$ and $N=100$.
3.2 Trust-as-reduced-monitoring across games
@@
We now investigate, as a function of the monitoring cost, (i) the frequency of the Trust-based Cooperation strategy TUC when all five strategies are possible, and (ii) the impact of adding trust-based strategies (TUC and TUD) to the pool of available strategies on the frequency of cooperation.
@@
Figure 4: Left: Frequency of strategies as a function of the monitoring cost, $\varepsilon$. Right: Frequency of cooperation in the absence or presence of trust-based strategies TUC and TUD, as a function of the monitoring cost, $\varepsilon$. The difference in frequency of cooperation when trust-based strategies are included is shaded in green where positive and red where negative. Results are presented for one example of each of the Prisoner’s Dilemma, Snowdrift and Stag-Hunt games. Parameters: trust-based strategies with $\theta=3$ and $p=0.25$, games with $r=50$, and evolutionary dynamics where $\beta=0.1$ and $N=100$.
@@
Figure 5: Frequency of the TUC strategy (left) and the impact of considering trust strategies on the frequency of cooperation (right), as a function of the temptation to defect $T$ and the sucker’s payoff $S$. The difference in frequency of cooperation is calculated between a system with all five strategies and a system with only AllC, AllD and TFT. The red dots correspond to the values of $S$ and $T$ used in Figure 4. Parameters: trust-based strategies with $\theta=3$ and $p=0.25$, games with $\varepsilon=0.25$ and $r=50$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
@@
3.2.1 Baseline analysis: Trust-based strategies are successful and promote cooperation in the repeated Prisoner’s Dilemma
@@
In the top part of Figure 4, which focuses on the PD, our analysis replicates previous findings [22]. First, increasing the monitoring cost diminishes the frequency of the TFT strategy (top left panel), since TFT must pay the monitoring cost every round. Second, a higher monitoring cost favours TUC compared to TFT, with TUC being the most successful strategy overall for intermediate monitoring costs in the PD. Third, the inclusion of trust-based strategies (TUC and TUD) in a system that previously included only AllC, AllD and TFT leads to an overall increase in the frequency of cooperation (top right panel). When the monitoring cost is high, the AllD strategy becomes the most prevalent, but (i) TUC remains among the most common across a wide range of monitoring costs, and (ii) the inclusion of trust-based strategies always increases the frequency of cooperation for the range of monitoring costs studied.
@@
3.2.2 Trust-based Cooperation (TUC) is a frequent strategy across different games
@@
Our setup now allows us to examine if these findings generalise across different types of games, and thus different types of social interaction. Our results in Figure 4 and the left panel of Figure 5 demonstrate that Trust-based Cooperation (TUC) consistently remains a frequent strategy across the three different games, and for a broad range of monitoring costs. In other words, trust-as-reduced-monitoring is expected to be a widespread heuristic across various types of interaction, not just in PD situations. A notable difference between the games is that trust-based strategies are highly prevalent in the PD, while AllC is almost absent. In this situation, cooperation with reduced monitoring after trust is established is a successful strategy, but unconditional cooperation is not. In the SD and SH games, both TUC and AllC strategies are frequent, indicating that a mix of careful trusters and unconditional cooperators could be observed in these more relaxed social dilemmas.
@@
3.2.3 The availability of trust-based strategies increases the frequency of cooperation when the temptation to defect is high
@@
Our results reveal that the impact of incorporating trust-based strategies on cooperation differs across the various games. Specifically, while the availability of the Trust-based Cooperation strategy is always beneficial in the PD, this strategy has a limited effect on the frequency of cooperation in SD games and in SH games. As shown in Figure 5, trust-based strategies generally increase cooperation when the temptation to defect (T) is high. However, trust-based strategies do not seem to increase cooperation in SH coordination games, even though TUC is present at a significant frequency. This is because in coordination games, we have shown earlier that reciprocal strategies are only a springboard for individuals that always cooperate to become established. Once in place, these unconditional strategies can enjoy a high payoff without ever paying a monitoring cost.
3.3 Trust-as-reduced-monitoring when individuals make errors
@@
Finally, we consider cases where agents can make unintentional errors in their actions. Figure 6 shows the frequency of the different strategies and the frequency of cooperation as a function of the error rate, $\mu_{\mathrm{e}}$. Our results show that an increasing error rate decreases the frequency of TUC across games, in favour of the strategy AllD. However, our results also show that the frequency of TUC remains considerable even for relatively large error rates such as 5%.
@@
Figure 6: Frequency of strategies (left) and frequency of cooperation in the absence or presence of trust-based strategies TUC and TUD (right), as a function of the error rate, $\mu_{\mathrm{e}}$. For clarity, the difference in frequency of cooperation is shaded in green when positive and red when negative. Parameters: trust-based strategies with $\theta=3$ and $p=0.25$, games with $\varepsilon=0.25$ and $r=50$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
@@
The right-hand side of Figure 6 shows that, in all games, error rates of up to 10% do not lessen the beneficial effect of trust on the frequency of cooperation, and can even increase it. For instance, TUC increases cooperation the most in PD games when the error rate is around 2%-3%. Importantly, trust-based strategies improve cooperation when the error rate is greater than 0 in SH games, in contrast to the results where agents do not make errors. A trust-as-reduced-monitoring heuristic thus helps agents achieve cooperation in coordination dilemmas where errors can occur.
@@
Trust-based strategies are thus selected for in the presence of errors, and can be even more beneficial for cooperation than in situations where agents never make errors. Figure 7 shows that these conclusions hold for games repeated for a larger number of rounds ($r=100$), and where trust-based strategies monitor less frequently once the trust threshold has been reached (high trustfulness, $p=0.04$).
@@
The effect of high trustfulness on cooperation varies with the type of game. Specifically, lowering $p$ from 0.25 to 0.04 increases cooperation in SD and SH games, but it decreases cooperation in the PD (right-hand panel of Figure 7). Thus, the optimum level of trustfulness for promoting cooperation depends on the type of social dilemma, with more relaxed dilemmas favouring greater trustfulness.
@@
Figure 7: Difference in the frequency of cooperation when trust-based strategies TUC and TUD are included, as a function of the error rate, $\mu_{\mathrm{e}}$, for three different social dilemmas and for long interactions or high trustfulness. Parameters: trust-based strategies with $\theta=3$, games with $\varepsilon=0.25$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
@@ 4 Discussion
Our results can be summarised in three points. First, we have shown that trust, as a heuristic to avoid the cost of always monitoring a partner’s actions, is a successful strategy that can evolve in a wide range of interactions. Importantly, this happens even in the presence of exploitative strategies (TUD) tailored to take advantage of this trust.
@@
Second, by disentangling behavioural measures of trust from behavioural measures of cooperation, we have been able to rigorously examine the effect that trust has on the overall level of cooperation in a population. Our results show that the presence of the Trust-based Cooperation strategy (TUC) increases the overall level of cooperation when the temptation to defect is high, but not in coordination dilemmas such as Stag-Hunt games. We have shown that this is because reciprocity in coordination games is useful to kickstart cooperation, but is not necessary once cooperation is established. This suggests that interventions aiming to promote the success of TUC, such as reducing high monitoring costs, would have less effect on cooperation in risky coordination games compared to stricter social dilemmas.
@@
Third, we have shown that when there is the possibility of agents making errors in their actions, TUC remains frequent and can even be more beneficial for cooperation compared to cases where agents never make errors. Notably, TUC does not suffer as much from errors as TFT, as TUC agents ignore errors from cooperative partners on the rounds that they do not monitor. By contrast, TFT is not trustful and monitors its partner’s actions every round, which means that a single mistake causes TFT to defect in retaliation. This can cause the partner to defect in retaliation, leading to a loss of cooperation over many rounds resulting from one mistake [45]. Trust can prevent this spiral into defection.
@@
Our model is based on the assumption that it can be costly to know whether the action of another agent was actually cooperative – even when you observe that action [22]. Everyday examples of this are common. For example, there is a time cost of monitoring if the correct amount has been debited from your bank account, or if goods that you have ordered online are genuine rather than fake. When interacting with artificial intelligence (AI), these costs can become even greater. For example, are personalised music recommendations from a streaming app influenced by paid promotions [46]? Has an LLM agent completed a task in a way that meets the preferences I prompted it with, or were its actions not in my interests due to biases in its training data [47]?
@@
Our operationalisation of trust – as choosing not to monitor another agent’s actions – provides an objective behavioural measure that applies across a broad spectrum of social interactions. Using this, trust can be measured in experimental settings by how often someone chooses (or not) to pay a cost to monitor another agent’s actions. This could be carried out in standard behavioural economics settings [48], where a participant is given the option to pay to find out the action of its partner in each round of a repeated social dilemma game. Some experimental studies have already considered costly monitoring for the particular case of the “trust game” [49, 50]. In this game, one player is the investor and the other player is the trustee. The investor begins with an endowment amount of money. They then choose a proportion (between 0 and 1 inclusive) of this to send to the trustee. The amount sent is multiplied by a factor >1>1. The trustee then chooses the proportion of this to return to the investor. When the game is repeated, most experiments assume that the investor is given complete information about the previous actions of the trustee at no cost, i.e. that monitoring is costless. However, a study has shown that when a cost is introduced to obtain this information, people are sensitive to this cost and monitor less [49]. Moreover, the frequency of monitoring was lower in later rounds, in a pattern akin to the TUC strategy. Future work should extend these experiments beyond the trust game and into the larger space of social dilemma games that we have modelled.
@@
Crucially, this trust measure can also be applied in more realistic experimental settings. One important application of this is experiments to measure the factors affecting people’s trust in AI. Although there have been many empirical studies in computer science that have aimed to measure factors affecting trust in various applications of AI, these studies have lacked a unified theoretical framework [51, 52]. This has led to conflicting results and a lack of predictive models, with “Trustworthiness mechanisms and measures […] being advanced in AI regulations and standards that may not actually increase trust” [53]. The trust measure we have introduced here can help to resolve this impasse, by providing an objective measure of trust that is grounded in (evolutionary) game theory. For example, incentivised experiments could be designed where a user can take advice from an AI agent in a realistic setting, such as an e-commerce chatbot recommending products or an LLM answering questions about a text, and can pay a cost to verify (monitor) whether that advice is correct. This can then be connected with EGT models by varying factors such as how frequently the AI gives correct information (i.e. its strategy to cooperate or defect), or whether the AI has a high reputation from other users.
@@
The costs of monitoring also lead to problems for the alignment and control of AI, that is, ensuring that the goals and behaviours of AI systems are aligned with human values and intentions. Research in AI safety has stressed the risk of scheming – AI systems that deceptively appear to be aligned with human goals during training, but then pursue misaligned goals after deployment once they are no longer closely monitored [54]. This type of scheming corresponds to the TUD strategy in our model – cooperate until an auditor or developer relaxes oversight, and then defect. The monitoring cost maps to the cost of auditing the AI system. Extensions of our model, with the payoff matrix calibrated to the parameters used in red-team blue-team AI control experiments [55], could therefore provide insights into auditing regimes and other institutional incentives that minimise vulnerability to scheming by LLMs and other future AI systems.
@@
Most generally, our model demonstrates the evolutionary success of a trust heuristic in competition with strategies that do not use trust (AllC, AllD and TFT), and in competition with strategies designed to build and then exploit trust (TUD). Trust evolves even though trust-based strategies are costly to use compared to unconditional strategies (since they pay the monitoring cost before trust is established and periodically thereafter), and are risky compared to standard reciprocal strategies (since they could potentially be exploited once they stop monitoring every round). These results suggest that a trust heuristic could have evolved in humans by natural selection. Indeed, there is evidence that human trust behaviour is to some extent genetically determined [56], and hence is an evolved trait. Our results suggest that repeated social dilemma games, as played by humans throughout (pre)history [57], could provide selection pressure for the evolution of trust.
@@
There are several limits to our model which could be addressed in future work. First, we have not modelled the evolution of the parameters of trust-based strategies that control when they monitor their co-player’s actions, $\theta$ and $p$. Rather, we have assumed that these are given exogenous values. Future work could allow these to be evolving traits, allowing evolutionary optimal values to be found for different game parameters.